# Random time-shift approximation enables hierarchical Bayesian inference of mechanistic within-host viral dynamics models on large datasets

### v1

@@ Abstract
Mechanistic mathematical models of within-host viral dynamics are tools for understanding how a virus’ biology and its interaction with the immune system shape the infectivity of a host. The biology of the process is encoded by the structure and parameters of the model that can be inferred statistically by fitting to viral load data. The main drawback of mechanistic models is that this inference is computationally expensive because the model must be repeatedly solved. This limits the size of the datasets that can be considered or the complexity of the models fitted. In this paper we develop a much cheaper inference method by implementing a novel approximation of the model dynamics that uses a combination of random and deterministic processes. This approximation also properly accounts for process noise early in the infection when cell and virion numbers are small, which is important for the viral dynamics but often overlooked. Our method runs on a consumer laptop and is fast enough to facilitate a full hierarchical Bayesian treatment of the problem with sharing of information to allow for individual level parameter differences. We apply our method to simulated data and a reanalysis of COVID-19 monitoring data in an National Basketball Association cohort of 163 individuals.
@@
Understanding how viruses reproduce within an infected host is crucial for predicting disease progression and evaluating treatments. One way to study this is by using mathematical models that describe how viruses reproduce in the body, such as in the respiratory tract. These models help understand key biological processes, but they require significant computing power, making it difficult to analyse data from large groups of individuals. In this study we developed a new, faster, method for analysing viral load data. Our approach combines randomness, to account for chance events early in infection, with a more predictable model once the infection is well established. This allows us to speed up calculations while capturing key biological processes and maintaining a high degree of accuracy. Our method is efficient enough to run on a laptop, making it possible to study large datasets using a full statistical approach that accounts for individual differences. We tested our method on simulated data and applied it to real-world data from 163 individuals in a COVID-19 monitoring study. Our results show that the model can accurately estimate key parameters and track viral load dynamics over time.
@@ 1 Introduction
Within-host viral dynamics (WHVD) models describe the progression of viral levels within an individual post-infection [1, 2, 3]. Depending on their complexity, these models can capture key biological processes such as viral replication, immune response, and clearance, or focus on summarising features like peak timing and growth/decline rates [1, 2, 3, 4]. In a typical acute infection, the amount of virus in an individual—referred to as the viral load (VL)—grows exponentially before peaking (due to the immune response and / or the carrying capacity of the respiratory tract) before declining [5]. Modelling this progression of virus inside a host can provide insight into important features like when an individual’s infectiousness (VL) peaks, differences between variants of a disease, and the vaccination status of individuals [6], all of which are drivers of between-host (or transmission) dynamics [7, 8, 9] and are key in informing health policies [10].
@@
Our second main contribution is that we can fit a fully stochastic model, instead of a only a deterministic one. To put this in context, mechanistic models are themselves broadly either stochastic or deterministic. Deterministic models are much cheaper to solve, but stochastic models capture random variation from when the populations are initially small (i.e. early in the infection) that can still have an impact on the time for the viral load to peak [16, 17]—ignoring this effect potentially biases inferences. The additional complexity of stochastic models has lead—with one exception [18]—to most inference approaches ignoring early time noise and only fitting simple deterministic models for the mean populations [11, 8, 5]. In this work we leverage the large scales over which the VL changes (from $10^{0}$ to $10^{8}$ virons per $\mu$l of plasma) to derive a novel approximation for the solutions of a fully stochastic WHVD model. This facilitates our first contribution by allowing the derivation of a fast and accurate approximation to the likelihood for VL time-series data. It also properly takes account of the early stochasticity in the process that is required for performing unbiased inference.
@@
Our method is so fast that it allows for fitting a full hierarchical Bayesian model to datasets of hundreds of individuals such as the COVID19 testing dataset collected from National Basketball Association (NBA) players during the 2020-2021 seasons [6, 4]. This dataset is one of the most complete and detailed datasets publicly available that features longitudinal measurements of individuals’ VL’s through time, using common data formats found in the literature [5, 11, 2, 19, 20]. In contrast to others [4, 6, 5, 18], we fit a full stochastic model to all 163 measured individuals and the inference algorithm is run on a consumer laptop in approximately one hour. The method can produce accurate posterior predictive VL trajectories and parameter estimates of the key within-host parameters (e.g. the within-host basic reproduction number, $R_{0}$) that are consistent with previous studies of respiratory illnesses [5, 11, 2, 19]. The hierarchical structure of the model allows us to share information from individuals with sufficient data to those whos data is limited, or noisy, during the growth and/or the decline phases of the infection. Our method can be easily extended to other datasets and other within-host models (i.e. models that include immune responses [8, 21]) and other viral infections and so provides a general tool for fitting to WHVD data.
@@
The key idea upon which our method builds is that the macroscopic dynamics of a fully stochastic WHVD model can be accurately approximated using a deterministic model with a random time-shift applied to the initial conditions [16, 17]. Fig. 1 illustrates this random time-shift idea. In Fig. 1A, the grey lines show realisations of a stochastic model (defined fully later) and the black line shows the deterministic approximation to the stochastic model. At early times the process is strongly affected by stochasticity due to the small numbers of cells and virions, but this occurs below a detection threshold (indicated by the horizontal red line) where data is not available. Once the virion numbers grow large enough, growth becomes exponential, and appears deterministic. The early time stochasticity is reflected on the macroscale (the scale on which measurements can be made, i.e. above the detection threshold) in the time for the virus to peak. We see that the deterministic model (black line) accurately captures the growth and decline rates as well as the peak height, but not the randomness in the time to peak. But, by time shifting the initial conditions of the deterministic trajectory—keeping the initial numbers fixed—the model can capture the randomness in the time to peak. The orange curve in Fig. 1B shows the density of this random variable; our previous work details how this can be accurately approximated by a closed form function that can be calculated for a given model [17].
{adjustwidth}
-1in0in Refer to caption
@@
Fig 1: Illustration of the time-shift concept. (A) Example simulations from the fully stochastic model (grey lines) and the deterministic approximation (black). The detection limit for observations is indicated by the red dashed line. (B) The same simulations as shown in panel (A) but starting the plot from when the simulations cross the detection limit. Translating the deterministic trajectory through time approximates the stochastic solutions on the macroscale. The orange curve shows the probability density of this shift, defined by aligning threshold-crossing times and is equivalent to a shift in the initial conditions of the deterministic model. The models and parameters are described in Sections 2.1 and 2.2, with $\boldsymbol{\theta}=(8,4,1.7,3,10,0)$ used in these simulations.
@@
Another way of looking at this approximation is that it allows for fast forward simulation of the model without the need for costly stochastic simulation, for example using the Gillespie algorithm [22, 23]. So for a given set of parameters, all that is required is a single solution of the deterministic model and samples from the corresponding time-shift distribution. In doing the inverse, i.e. inference of model parameters from data, adopting the time-shift approximation is also enlightening. In many existing studies of this problem the initial viral load is treated as being unobserved and hence a parameter to be inferred [11] or fixed at some value [5]. The time-shift idea facilitates a shift in perspective from inferring the initial viral load at a fixed time, to instead inferring a fixed viral load at a random time. This turns out to be a key insight as it allows the derivation of a number of further accurate approximations for the likelihood function that provide large computational speed ups, allowing inference on large datasets using a relatively complicated model.
We begin in Section 2 by outlining our particular stochastic viral kinetics model and show mathematically how the dynamics can be approximated using the time-shift ideas. We then provide details of approximations that can be made as a result of how the time-shift is incorporated into the likelihood in order to improve the efficiency of our inference method. This provides the foundations for a Metropolis-Hastings within Gibbs sampler for sampling from the posterior distribution. In Section 3 we apply it to two datasets. Firstly, a synthetic data set that is used to validate the method and quantify its inference capabilities in a context where the true values of the parameters are known. We find that our approach is able to disentangle the measurement noise from the process noise and results in parameter estimates that are more consistent with the simulated parameters. Secondly, we fit the model to the NBA COVID-19 dataset [6, 4] where we find that our posterior predictive fits indicate that the method is able to produce a realistic fit to the data and obtain $R_{0}$ estimates that are consistent with previous studies. Additionally we show that the method is able to share information across individuals, to those with few observations and missing data, to estimate their viral dynamics. All the data and code are available in the following GitHub repository: https://github.com/djmorris7/random_time_shift-within_host_inference.
@@ 2.1 The viral kinetics model
In this work we consider one of the simplest viral kinetics model referred to as the Target-Cell-Limited ($\mathcal{TCL}$) model with an eclipse-phase [2, 24, 25]. This model captures the basic cell infection cycle and viral production cycles observed during an infection while accounting for an appropriate delay (the eclipse phase) between a cell becoming infected and producing virus. We choose this simpler model and not something more complex (see [14] for an example) as some of the parameters are already unidentifiable from VL data [26, 25] and so adding more complexity without additional data sources only exacerbates this. The method itself can easily be applied to more complex models if required. The $\mathcal{TCL}$ model is commonly specified as a set of (deterministic) ordinary differential equations [2, 24, 27] for the total population counts, but here we will instead start with a stochastic version that captures the individual nature of the populations from which we will derive the random time-shift approximation consisting of a deterministic ODE and the time-shift distribution itself.
@@
Infection (1,+1, 0, 0)11 0 0(-1,\ +1,\ 0,\ 0)( - 1 , + 1 , 0 , 0 ) $\beta S(t)V(t)$
Cell becomes infectious (0,1,+1, 0)011 0(0,\ -1,\ +1,\ 0)( 0 , - 1 , + 1 , 0 ) $kE(t)$
Infectious cell recovery (0, 0,1, 0)0 01 0(0,\ 0,\ -1,\ 0)( 0 , 0 , - 1 , 0 ) $\delta I(t)$
Production of virus (0, 0, 0,+1)0 0 01(0,\ 0,\ 0,\ +1)( 0 , 0 , 0 , + 1 ) $\rho V(t)$
Clearance of virus (0, 0, 0,1)0 0 01(0,\ 0,\ 0,\ -1)( 0 , 0 , 0 , - 1 ) $cV(t)$
@@
There is a choice to be made regarding the initial condition, $\boldsymbol{X}(0)$. The infection is triggered by the entry of virus into the system, so ideally, the initial viral load, $V(0)$, would be set. However, this introduces complexity regarding the amount of virus required to cause an infection [12, 31]. Jones et al. [31] show that the dynamics of the model are not significantly altered by taking the initial condition as a single infected cell, as any initial viral particles must initiate an infection by infecting one or more cells (at a rapid rate) before being cleared. Throughout this work we assume that $S(0)=8\times 10^{7}$ based on Zitzmann et al. [5]. Hence, throughout this work, we consider viral infections triggered by a single infected (but not actively infectious) cell, i.e., the initial condition is $\boldsymbol{X}(0)=(S(0)-1,1,0,0)$. The methods presented here can be extended to handle more complex initial conditions, and extensions of this are discussed in Section 4.
@@ 2.2 Random time-shift approximation
Here, we show that the macroscopic dynamics of the stochastic model described in Section 2.1 can be approximated by a deterministic ODE model, which corresponds to standard formulations with an added random time shift to the initial conditions. The deterministic approximation follows from a mean-field approximation of the CTMC model in the limit $K\to\infty$ (i.e., for large volume). In the context of respiratory infections, $K$ corresponds to the total volume over which the within-host infection process occurs and can therefore be considered “large” in the sense that the number of interacting particles (e.g., cells and virions) is sufficiently high. Throughout this work, however, the exact value of $K$ is unimportant, as it can be absorbed into the model parameters. Let $s(t)=S(t)/K$, $e(t)=E(t)/K$, $i(t)=I(t)/K$ and $v(t)=V(t)/K$ denote the densities of the cell and virus populations of each type, then in the limit as $K\to\infty$ these evolve via a set of coupled ODEs (see Section 1 of S1 text for derivation) [35])
[…]
We can multiply Eqs. (2) through by the volume $K$ as it is large and fixed to obtain the system of ODEs that govern the approximation for the mean dynamics [5, 2]. We use the notation $S_{d}(t)=Ks(t),E_{d}(t)=Ke(t)$, etc. with the subscript $d$ identifying these variables as solutions of the deterministic system. Once more, for consistency between the stochastic and deterministic formulations, we make the substitution $\beta=\alpha/K$,
[…]
@@
These equations match the ODEs that are usually just written down when working in a purely deterministic modelling paradigm [5, 3]. It is difficult to work with the transmission parameter $\beta$ as it is strongly dependent on the experimental setup in the study of interest (i.e. what method is used to collect VL samples) and negatively correlated with $\rho$ [5] so in this work we reparameterise the model in terms of $R_{0}$ (and back transform to obtain $\beta$ when needed) and hence the parameters are collected in the vector $\boldsymbol{\theta}=(R_{0},k,\delta,\rho,c,t_{0}$).
@@
Denote the solution of Eqs. (3), $\boldsymbol{X}_{d}(t)=(S_{d}(t),E_{d}(t),I_{d}(t),V_{d}(t))$, then this provides an approximation to the mean dynamics of the full stochastic CTMC $\{\boldsymbol{X}(t),t\geq t_{0}\}$, but does not capture the effect of the noise during the early-time phase when the numbers of infected cells and virions is low (see Fig. 1 and discussion in Section 1). The predominant effect of the stochasticity is to shift the time at which the peak VL occurs and subsequently the time at which the infection is cleared [16, 36, 17]. The work of Barbour et al. [16] formalises the connection between the shifted mean field kinetics and samples of the full stochastic process. They show that $\{\boldsymbol{X}(t),t\geq t_{0}\}$ is well approximated on the macroscale, i.e. the growth and decline rates and peak time and height, by another stochastic process $\{\boldsymbol{X}_{d}(t+\tau),t\geq t_{0}\}$ where $\tau:=\tau(\boldsymbol{\theta})\in\{-\infty\}\cup\mathbb{R}$ is a scalar random variable referred to as the random time-shift [16, 17].
@@
The practical importance of this result is that sample paths of $\{\boldsymbol{X}(t),t\geq t_{0}\}$ resemble those of the deterministic model, $\boldsymbol{X}_{d}(t)$, except for a translation in time by the time-shift $\tau$. The case $\tau=-\infty$ corresponds to when the number of virions and infected cells drops to zero, i.e. extinction of the virus within the host. Our interest lies in the modelled VL once $V(t)$ has grown to be measurable, i.e. outside of the early-time small population regime. Approximating the stochastic process $\{\boldsymbol{X}(t),t\geq t_{0}\}$ in this way means we can efficiently generate sample paths $V(t)$ by solving for the mean field kinetics and obtaining the distribution of the time-shift. The process $\{\boldsymbol{X}_{d}(t+\tau),t\geq t_{0}\}$ does not capture the noisy early dynamics (i.e. the actual stochastic fluctuations in the populations when these counts are small) but this is of little consequence as small VL’s fall below a detection threshold and are therefore not measurable [5, 4, 11].
The distribution of the random variable $\tau$ is derived by analysing a branching process approximation to the early time dynamics and can be expressed as the mixture [17]
[…]
where $q^{\star}(\boldsymbol{\theta})$ is the probability of ultimate extinction and the term inside the integrand, $f(\tau\,|\,\boldsymbol{\theta})$, is the density function for $\tau$ conditional on non-extinction. As a result of the simplicity of the $\mathcal{TCL}$ model, $q^{\star}(\boldsymbol{\theta})$, can be solved for analytically by solving a set of non-linear equations (see Section 2 of S1 text). In previous work [17] we showed that the density function for $\tau$, conditional on non-extinction, can be accurately approximated by a member of a family of functions with the form
@@
with $(a,d,p)$ determining the particular member of the family. The quantity $\lambda$—the early growth rate of the model—and the time-shift parameters $\boldsymbol{\zeta}=(a,p,d)$ are numerically calculated from the branching process approximation to the CTMC (see [17] for concrete examples). The red curve shown in Fig. 1 is this density evaluated for the given parameters. Recall that as $\tau$ is a function of $\boldsymbol{\theta}$, the parameters $\boldsymbol{\zeta}$ are also a function of $\boldsymbol{\theta}$ but that this has been suppressed in the notation for clarity. One final and important point is that the time-shift is independent of the infection time $t_{0}$.

### v2

@@ Abstract
Mechanistic mathematical models of within-host viral dynamics are tools for understanding how a virus’ biology and its interaction with the immune system shape the infectivity of a host. The biology of the process is encoded by the structure and parameters of the model that can be inferred statistically by fitting to viral load data. The main drawback of mechanistic models is that this inference is computationally expensive because the model must be repeatedly solved. This limits the size of the datasets that can be considered or the complexity of the models fitted. In this paper we develop a much cheaper inference method by implementing a novel approximation of the model dynamics that uses a combination of random and deterministic processes. This approximation also properly accounts for process noise early in the infection when cell and virion numbers are small, which is important for the viral dynamics but often overlooked. Our method runs on a consumer laptop and is fast enough to facilitate a full hierarchical Bayesian treatment of the problem with sharing of information to allow for individual level parameter differences. We apply our method to simulated datasets and a reanalysis of COVID-19 monitoring data in an National Basketball Association cohort of 163 individuals.
@@
Understanding how viruses reproduce within an infected host is crucial for predicting disease progression and evaluating treatments. One way to study this is by using mathematical models that describe how viruses reproduce in the body, such as in the respiratory tract. These models help understand key biological processes, but they require significant computing power, making it difficult to analyse data from large groups of individuals. In this study we developed a new, faster, method for analysing viral load data. Our approach combines randomness, to account for chance events early in infection, with a more predictable model once the infection is well established. This allows us to speed up calculations while capturing key biological processes and maintaining a high degree of accuracy. Our method is efficient enough to run on a laptop, making it possible to study large datasets using a full statistical approach that accounts for individual differences. We tested our method on 50 simulated datasets and applied it to real-world data from 163 individuals in a COVID-19 monitoring study. Our results show that the model can accurately estimate key parameters and track viral load dynamics over time.
@@ 1 Introduction
Within-host viral dynamics (WHVD) models describe the progression of viral levels within an individual post-infection [1, 2, 3]. Depending on their complexity, these models can capture key biological processes such as viral replication, immune response, and clearance, or focus on summarising features like peak timing and growth/decline rates [1, 2, 3, 4]. In a typical acute infection, the amount of virus in an individual—referred to as the viral load (VL)—grows exponentially before peaking (due to the immune response and/or the carrying capacity of the respiratory tract) and then declining [5]. Modelling this progression of virus inside a host can provide insight into important features like when an individual’s infectiousness (VL) peaks, differences between variants of a disease, and the vaccination status of individuals [6], all of which are drivers of between-host (or transmission) dynamics [7, 8, 9] and are key in informing health policies [10].
@@
[…]
@@
Our method is fast enough to fit a full hierarchical Bayesian model to datasets of hundreds of individuals, such as the COVID-19 testing dataset collected from National Basketball Association (NBA) players during the 2020–2021 seasons [6, 4]. This dataset is among the most complete and detailed publicly available, featuring longitudinal measurements of individuals’ VLs over time using common data formats found in the literature [5, 11, 2, 19, 20]. In contrast to others [4, 6, 5, 18], we fit a full stochastic model to the 163 measured individuals and the inference algorithm is run on a consumer laptop in approximately one hour. The method can produce accurate posterior predictive VL trajectories and parameter estimates of the key within-host parameters (e.g. the within-host basic reproduction number, $R_{0}$) that are consistent with previous studies of respiratory illnesses [5, 11, 2, 19]. The hierarchical structure of the model allows us to share information from individuals with sufficient data to those whose data is limited, or noisy, during the growth and/or the decline phases of the infection. Our method can be easily extended to other datasets and other within-host models (i.e. models that include immune responses [8, 21]) and other viral infections and so provides a general tool for fitting to WHVD data.
@@
The key idea upon which our method builds is that the macroscopic dynamics of a fully stochastic WHVD model can be accurately approximated using a deterministic model with a random time-shift applied to the initial conditions [16, 17]. Fig. 1 illustrates this random time-shift idea. In Fig. 1A, the grey lines show realisations of a stochastic model (defined fully later) and the black line shows the deterministic approximation to the stochastic model. At early times the process is strongly affected by stochasticity due to the small numbers of cells and virions, but this occurs below a detection threshold (indicated by the horizontal red line) where data is not available. Once the virion numbers grow large enough, growth becomes exponential, and appears deterministic. The early time stochasticity is reflected on the macroscale (the scale on which measurements can be made, i.e. above the detection threshold) in the time for the virus to peak. We see that the deterministic model (black line) accurately captures the growth and decline rates as well as the peak height, but not the randomness in the time to peak. But, by time shifting the initial conditions of the deterministic trajectory—keeping the initial numbers fixed—the model can capture the randomness in the time to peak. The blue curve in Fig. 1B shows the density of this random variable; our previous work details how this can be accurately approximated by a closed form function that can be calculated for a given model [17].
@@
Fig 1: Illustration of the time-shift concept. (A) Example simulations from the fully stochastic model (grey lines) and the deterministic approximation (black). The detection limit for observations is indicated by the red dashed line. (B) The same simulations as shown in panel (A) but starting the plot from when the simulations cross the detection limit. Translating the deterministic trajectory through time approximates the stochastic solutions on the macroscale. The blue curve shows the probability density of this shift, defined by aligning threshold-crossing times and is equivalent to a shift in the initial conditions of the deterministic model. The models and parameters are described in Sections 2.1 and 2.2, with $\boldsymbol{\theta}=(8,4,1.7,3,10,0)$ used in these simulations.
@@
Another way of looking at this approximation is that it allows fast forward simulation of the model without the need for costly stochastic simulation, for example using the Gillespie algorithm [22, 23]. For any given set of parameters, only a single solution of the deterministic model and samples from the corresponding time-shift distribution are required. In many existing studies of this problem the initial viral load is treated as being unobserved and hence a parameter to be inferred [11] or fixed at some value [5]. The time-shift idea facilitates a shift in perspective from inferring the initial variable viral load at a fixed time, to instead inferring a fixed viral load at a random time. This turns out to be a key insight as it allows the derivation of a number of further accurate approximations for the likelihood function that provide large computational speed ups, allowing inference on large datasets using a relatively complicated model.
We begin in Section 2 by outlining our particular stochastic viral kinetics model and show mathematically how the dynamics can be approximated using the time-shift ideas. We then provide details of approximations that can be made as a result of how the time-shift is incorporated into the likelihood in order to improve the efficiency of our inference method. This provides the foundations for a Metropolis-Hastings within Gibbs sampler for sampling from the posterior distribution. In Section 3 we apply it to two types of datasets. Firstly, multiple synthetic datasets are used to validate the method and quantify its inference capabilities in contexts where the true values of the parameters are known. We find that our approach is able to disentangle the measurement noise from the process noise and results in parameter estimates that are consistent with the simulated parameters. Secondly, we fit the model to the NBA COVID-19 dataset [6, 4], where we find that our posterior predictive fits indicate that the method is able to produce a realistic fit to the data and obtain $R_{0}$ estimates that are consistent with previous studies. Additionally, we show that the method is able to share information across individuals, to those with few observations and missing data, to estimate their viral dynamics. All the data and code are available in the following GitHub repository: https://github.com/djmorris7/random_time_shift-within_host_inference.
@@ 2.1 The viral kinetics model
In this work we consider one of the simplest viral kinetics model referred to as the Target-Cell-Limited ($\mathcal{TCL}$) model with an eclipse-phase [2, 24, 25]. This model captures the basic cell infection cycle and viral production cycles observed during an infection while accounting for an appropriate delay (the eclipse phase) between a cell becoming infected and producing virus. We choose this simpler model and not something more complex (see [14] for an example) as some of the parameters are already unidentifiable from VL data [26, 25] and so adding more complexity without additional data sources only exacerbates this. The method itself can easily be applied to more complex models if required. The $\mathcal{TCL}$ model is commonly specified as a set of deterministic ordinary differential equations [2, 24, 27] for the total population counts, but here we will instead start with a stochastic version that captures the individual nature of the populations from which we will derive the random time-shift approximation consisting of a deterministic ODE and the time-shift distribution itself.
@@
Infection (1,+1, 0, 0)(-1,\ +1,\ 0,\ 0) $\beta S(t)V(t)$
Cell becomes infectious (0,1,+1, 0)(0,\ -1,\ +1,\ 0) $kE(t)$
Infectious cell recovery (0, 0,1, 0)(0,\ 0,\ -1,\ 0) $\delta I(t)$
Production of virus (0, 0, 0,+1)(0,\ 0,\ 0,\ +1) $\rho V(t)$
Clearance of virus (0, 0, 0,1)(0,\ 0,\ 0,\ -1) $cV(t)$
@@
There is a choice to be made regarding the initial condition, $\boldsymbol{X}(0)$. The infection is triggered by the entry of virus into the system, so ideally, the initial viral load, $V(0)$, would be set. However, this introduces complexity regarding the amount of virus required to cause an infection [12, 31]. Smith et al. [31] show that the dynamics of the model are not significantly altered by taking the initial condition as a single infected cell, as any initial viral particles must initiate an infection by infecting one or more cells (at a rapid rate) before being cleared. Throughout this work we assume that $S(0)=8\times 10^{7}$ based on Zitzmann et al. [5]. Hence, throughout this work, we consider viral infections triggered by a single infected (but not actively infectious) cell, i.e., the initial condition is $\boldsymbol{X}(0)=(S(0)-1,1,0,0)$. The methods presented here can be extended to handle more complex initial conditions, and extensions of this are discussed in Section 4.
@@ 2.2 Random time-shift approximation
Here, we show that the macroscopic dynamics of the stochastic model described in Section 2.1 can be approximated by a deterministic ODE model, which corresponds to standard formulations with an added random time shift to the initial conditions. The deterministic approximation follows from a mean-field approximation of the CTMC model in the limit $K\to\infty$ (i.e., for large volume). In the context of respiratory infections, $K$ corresponds to the total volume over which the within-host infection process occurs and can therefore be considered “large” in the sense that the number of interacting particles (e.g., cells and virions) is sufficiently high. Throughout this work, however, the exact value of $K$ is unimportant, as it can be absorbed into the model parameters. Taking the large-volume limit yields a system of coupled ODEs for the cell and virus populations (see Section 1 of S1 text. for the derivation). This system governs the mean-field dynamics, denoted $S_{d}(t),E_{d}(t),I_{d}(t),V_{d}(t)$, where the subscript $d$ indicates deterministic solutions. To maintain consistency between the stochastic and deterministic formulations, we again make the substitution $\beta=\alpha/K$, giving
$\begin{split}\frac{\operatorname{d\!}{S_{d}(t)}}{\operatorname{d\!}{t}}&=-\beta S_{d}(t)V_{d}(t),\\ \frac{\operatorname{d\!}{E_{d}(t)}}{\operatorname{d\!}{t}}&=\beta S_{d}(t)V_{d}(t)-kE_{d}(t),\\ \frac{\operatorname{d\!}{I_{d}(t)}}{\operatorname{d\!}{t}}&=kE_{d}(t)-\delta I_{d}(t),\\ \frac{\operatorname{d\!}{V_{d}(t)}}{\operatorname{d\!}{t}}&=\rho I_{d}(t)-cV_{d}(t).\end{split}$ (2)
These equations coincide with the ODEs typically used in purely deterministic modelling approaches [5, 3].
@@
It is difficult to work directly with the transmission parameter $\beta$ as it is strongly dependent on the experimental setup in the study of interest (e.g., how viral load samples are collected) and negatively correlated with $\rho$ [5]. Therefore, we reparameterise the model in terms of $R_{0}$ and back-transform to obtain $\beta$ when needed. We collect all the parameters in the vector $\boldsymbol{\theta}=(R_{0},k,\delta,\rho,c,t_{0})$.
@@
Denote the solution of Eqs. (2), $\boldsymbol{X}_{d}(t)=(S_{d}(t),E_{d}(t),I_{d}(t),V_{d}(t))$, then this provides an approximation to the mean dynamics of the full stochastic CTMC $\{\boldsymbol{X}(t),t\geq t_{0}\}$, but does not capture the effect of the noise during the early-time phase when the numbers of infected cells and virions is low (see Fig. 1 and discussion in Section 1). The predominant effect of the stochasticity is to shift the time at which the peak VL occurs and subsequently the time at which the infection is cleared [16, 35, 17]. The work of Barbour et al. [16] formalises the connection between the shifted mean field kinetics and samples of the full stochastic process. They show that $\{\boldsymbol{X}(t),t\geq t_{0}\}$ is well approximated on the macroscale, i.e. the growth and decline rates and peak time and height, by another stochastic process $\{\boldsymbol{X}_{d}(t+\tau),t\geq t_{0}\}$ where $\tau\mathrel{\mathop{\ordinarycolon}}=\tau(\boldsymbol{\theta})\in\{-\infty\}\cup\mathbb{R}$ is a scalar random variable referred to as the random time-shift [16, 17].
@@
The practical importance of this result is that sample paths of $\{\boldsymbol{X}(t),t\geq t_{0}\}$ resemble those of the deterministic model, $\boldsymbol{X}_{d}(t)$, except for a translation in time by the time-shift $\tau$. The case $\tau=$-$\infty$ corresponds to when the number of virions and infected cells drops to zero, i.e. extinction of the virus within the host. Our interest lies in the modelled VL once $V(t)$ has grown to be measurable, i.e. outside of the early-time small population regime. Approximating the stochastic process $\{\boldsymbol{X}(t),t\geq t_{0}\}$ in this way means we can efficiently generate sample paths $V(t)$ by solving for the mean field kinetics and obtaining the distribution of the time-shift. The process $\{\boldsymbol{X}_{d}(t+\tau),t\geq t_{0}\}$ does not capture the noisy early dynamics (i.e. the actual stochastic fluctuations in the populations when these counts are small) but this is of little consequence as small VLs fall below a detection threshold and are therefore not measurable [5, 4, 11].
The distribution of the random variable $\tau$ is derived by analysing a branching process approximation to the early time dynamics [17]. Previous work [17] showed that the density function for $\tau$, conditional on non-extinction, can be accurately approximated by a member of a family of functions with the form
@@
with $(a,d,p)$ determining the particular member of the family. The quantity $\lambda$—the early growth rate of the model—and the time-shift parameters $\boldsymbol{\zeta}=(a,p,d)$ are numerically calculated from the branching process approximation to the CTMC (see [17] for concrete examples). The blue curve shown in Fig. 1 is this density evaluated for the parameters given in the caption. Recall that as $\tau$ is a function of $\boldsymbol{\theta}$, the parameters $\boldsymbol{\zeta}$ are also a function of $\boldsymbol{\theta}$ but that this has been suppressed in the notation for clarity. One final and important point is that the time-shift is independent of the infection time $t_{0}$.
Computing the parameters of a single time-shift distribution is not expensive, but when incorporated into an Markov chain Monte Carlo (MCMC) routine, where they must be repeatedly evaluated, it can become a bottleneck. To overcome this, we train a neural network (NN) to learn the function $g\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{5}\mapsto\mathbb{R}^{3}$, mapping $\boldsymbol{\theta}=(R_{0},k,\delta,\rho,c)$ to $\boldsymbol{\zeta}=(a,b,p)$. This approach, known as amortised optimisation [36, 37, 38], replaces repeated optimisation with a learned mapping, significantly decreasing the cost of computation during the running of the algorithm, but at the expense of initially training the network [38]. Full details of our implementation are provided in Section 6 of S1 text.. In testing we found that this modification to our method accelerates the computation of the time-shift parameters by roughly 200 times.